UESTION 32 An income statement for Don & Dar Company is presented below. Don & Dar Company Income Statement For the Year Ended January 31, 2018 1,800,000 Sales (300,000 units) Variable Expenses $900,000 225,000 Cost of Goods Sold Administrative Expense 150,000 Fixed Costs: 1,275,000 525,000 Contribution Margin Cost of Goods Sold 400,000 Selling Expenses Administrative Expense 100,000 242,400 742.400 钆217,400) Net Loss INSTRUCTIONS (MUST SHOW ALL COMPUTATIONS) 1 Compute the break -even in total sales dollars and in units for 2018 has proposed a plan to get the enterpise on a profitable level He proposes to improve the quality of the product by spending S0.30 more on raw materials The seling price persrit would be increased by $0.35 The sales volume would ncrease by 25% How wouldDons 3 Dar want to improve profits vis marketing s smoooo. Man n etarom tar belie and promotionsHis plan is to increase seling expense by S0.15 lower the selling price by S0.35 per unit, slae sales volume would increase by 60% if thesertha sel rar price by S 35 per unit in nee ei l s e a, How wodd Dars proposal would increase by 60% if these changes are made. 4 Which plan do you recommend and why?
Answer #1

1. Breakeven point is the activity level at which there is neither profit nor loss.

   Breakeven point(units)= Total fixed costs \tiny \div contibution per unit

      Contribution per unit= Total contribution \tiny \div Number of units

=$525,000\tiny \div300,000 units

   = $1.75/unit

Therefore,Breakeven point in terms of units=$742,400(Total Fixed Costs)\tiny \div 1.75 (contribution per unit)

=424,229 units,ie,this many units need to be sold in order for the company to break even

Breakeven point(sales)= Total Fixed costs\tiny \div Contribution/sales ratio

        The contribution/sales ratio= contribution\tiny \divsales

                                                =$525,000\tiny \div$1,800,000

                                                 =29%             

Therefore,Breakeven point in terms of sales value=$742,400\tiny \div0.29

                                             =$2,545,430,ie,This amount worth of sales need to happen for the company to break even

2. At this point,I think it'd be helpful to set out the per unit costs and price of the product.

Selling price/unit=total Sales value/number of units sold

                       = 1800,000/300,000

                        =$6/unit

The Variable cost/unit has to be split into its constituent parts:

  •     COGS/unit= Total COGS/number of units

                               =900,000/300,000

                                =$3/unit

  • Similarly,Selling expense/unit=$.75/unit
  • Administrative expense/unit=$0.5/unit

The adjustments that Don wishes to make are:

  • An increase of $0.3 in raw material cost. Therefore,COGS=3+0.3=$3.3/unit
  • An increase of $.035 in selling price.Therefore,SP=6+0.35=$6.35/unit
  • An increase in sales volume of 25%.Sales volume=300,000+25%=375,000 units

The Income Statement would then resemble:

Sales($6.35*375,000 units) 2,381,250
Variable costs:
COGS($3.3*375,000 units) 1,237,500
Selling expense($0.75*375,000 units) 281,250
Administrative expense($0.5*375,000 units) 187,500 (1,706,250)
Contribution margin 675,000
Fixed costs( doesn't change) (742,400)
Net loss (67,400)

The net loss decreases by $150,000(217,400-67,400).This plan, although it still doesn't achieve a profit,does earn much less net loss that before.

The new breakeven point in sales would be= 742,400\tiny \div 0.283 [see below]

                                                               =$2,623,322

If this plan is followed through,The breakeven sales would increase by $77,892 (2,623,322-2,545,430),ie,the company would have to sell that much amount worth of products more for it to now break even.

The C/S ratio=675,000/2,381,250

   =28.3%   

3. Dar's plan would require the following adjustments:

  • An increase of $0.15 in selling expense.Therefore, selling expense=0.75+0.15=$0.9/unit
  • A decrease of $0.35 in Selling price.SP=6-0.35=$5.65/unit
  • An increase of $40,000 in fixed selling expenses.Fixed selling expense=242,400+40,000=$282,400
  • An increase of 60% in sales volume.Sales volume=300,000+60%=480,000 units

If we proceed with Dar's proposal,the income statement would look like:

Sales($5.65*480,000 units) 2,712,000
Variable expense:
COGS($3*480,000) 1,440,000
Selling expense($0.9*480,000) 432,000
Administrative expense($0.5*480,000 units) 240,000 (2,112,000)
Contribution margin 600,000
Fixed costs:
COGS 400,000
Selling expense 282,400
Administrative expense 100,000 (782,400)
Net loss (182,400)

   The net loss would decrease by an amount of $35,000( 217,400-182,400).

C/S ratio=600,000/2,712,000

=22%

Therefore,Breakeven point in terms of sales=782,400/0.22

                                                              =$3,556,364

The company would have to sell $1,010,934 (3,556,364-2,545,430)worth of products more to breakeven,if Dar's proposal is chosen

d) Comparing the 2 proposals:

DON'S PLAN DAR'S PLAN
Net loss 67,400 182,400
C/S ratio 28.3% 22%
Breakeven sales 2,623,322 3,556,364
  • Net loss: This is pretty obvious.Losses should always be as low as possible,if not completely avoidable.Therefore,the better option would be the Don's plan,since it offers a lesser amount of loss.
  • C/S ratio: c/s ratio should always be as high as possible.It gives a measure of how much contribution each $ of sale makes.In this case,Don's proposal achieves the higher C/S ratio.
  • Breakeven sales: This value shows the amount of sales that needs to be acheived before the company breaks even.The lower this value,the better.Don's proposal offers a much lower breakeven sales amount.

In all 3 criterias mentioned above,Don's plan seems more likely to acheive better results than Dar's.For all the aforementioned reasons,I'd recommend Don's proposal.